I HAVE A TWO DIGIT NUMBER. THE SUM OF 4 TIMES THE TENS DIGIT AND THE UNITS DIGIT IS 26. THE SUM OF THE TENS DIGIT AND 4 TIMES THE UNITS DIGIT IS 14. WHAT IS THE SUM OF THE DIGITS OF MY NUMBER?
Answer by Anthea Lawn(22)

Let t be the tens digit and u be the units digit.
In that case, the number can be written as 10t + u (eg 43 is 4 X 10 + 3)
We are told :
4t + u = 26 -------(1)
And we are also told :
t + 4u = 14 ----------(2)
multiplying (2) by 4 gives :
4t + 16u = 56 ------(3)
and subtracting (1) gives :
15u = 30
u=2
therefore t=6
Hence the number is 62 and the sum of its digits is 8.